OpenStudy (anonymous):

Solve: −4(x + 1) − 3 = −3(x − 4) −19 19 2 0

OpenStudy (anonymous):

what grade r u in

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

lol and you get these easy questions

OpenStudy (anonymous):

yeah, but idk how to do it... -_-

OpenStudy (anonymous):

it is -19

OpenStudy (anonymous):

yeah its -19

OpenStudy (anonymous):

K so we have to distribute -4(x+1) and on the other side we got to distribute -3(x-4)   Now we have- -4x-7=-3x+12. We add 3x on both sides and that gives us +3x +3x -------------- -x-7=12 [Now we add 7 to both sides] +7 +7 --------------- -x=19 [Finally we multiply both sides by -1 so that the x can be positive] x=-19 -Hope that helps you understand it better :)